An unknown nonelectrolyte is dissolved in 1kg of solvent and the boiling point is increased by 10oC. Given that the boiling point elevation constant (Kb) is 1.00oC/m, determine the number of moles of solute present in the solution.

a. 1 mol
b. 5 mol
c. 8 mol
d. 10 mol
1
Expert's answer
2020-04-22T09:19:52-0400

Solution.

"\Delta T = Kb \times Cm"

Cm = 10 mol/kg

"Cm = \frac{n}{m(solvent)}"

"n = Cm \times m(solvent)"

n = 10 mol

Answer:

d